### v1

# A formal implementation of Behavior Trees to act in robotics

Abstract
Behavior Trees (BT) are becoming quite popular as an Acting component of autonomous robotic systems. We propose to define a formal semantics to BT by translating them to a formal language which enables us to perform verification of programs written with BT, as well as runtime verification while these BT execute. This allows us to formally verify BT correctness without requiring BT programmers to master formal language and without compromising BT most valuable features: modularity, flexibility and reusability. We present the formal framework we use: Fiacre, its langage and the produced TTS model; Tina, its model checking tools and Hippo, its runtime verification engine. We then show how the translation from BT to Fiacre is automatically done, the type of formal LTL and CTL properties we can check offline and how to execute the formal model online in place of a regular BT engine. We illustrate our approach on two robotics applications, and show how BT could benefit of other features available in the Fiacre formal framework (state variables, time, etc).
1 Introduction and motivation
Behavior Trees (BT) were initially developed and deployed to program Non-Player Characters (NPCs) in video games. BT are a powerful and modular framework for designing and implementing decision-making and control systems. They provide a structured way to define complex behaviors by combining smaller, reusable behavior modules into a hierarchical tree structure. BT have gained popularity in robotics due to their flexibility, readability, and ability to handle dynamic environments (e.g., Nav2, a popular navigation stack used in ROS2, is now deployed using BT Macenski et al. (2024), similarly, BT are a keystone of projects such as Street et al. (2024) where they aim at verifying the whole development toolchain).
Why yet another BT implementation? To provide a formal representation of BT which can then be used to formally verify interesting properties of the BT (can this node succeed, or fail, is this node reachable or not, can we prove that if we reach this state, eventually we will reach this one within some time interval, etc), and also to execute the formal model which implements the BT in lieu of a regular BT engine.
A behavior tree consists of nodes arranged hierarchically. Parent nodes tick their children to pass them the execution “token”. Children return either success, running or failure to their parent when the current execution tick is done, and the parent pursues the execution following a specified behavior.
Root node
The entry point of the tree that initiates the behavior evaluation and generates the successive ticks.
Leaf nodes
The terminal nodes that perform specific functions when ticked:
Condition nodes
Evaluate conditions, such as checking sensor data, battery level, or environmental states. They return success or failure.
Action nodes
Trigger actions in the environment, such as moving a joint, picking up an object, or sending a command. Action, on top of success or failure, may also return running (e.g., when it takes some time to complete the action such as a robot motion).
Control nodes
These modes specify the execution logic (behavior) of their children nodes (i.e., how to execute them):
Sequence
Executes its children from left to right. If any child fails, the Sequence fails, halting further execution. If the last one succeeds, the Sequence succeeds. If a child returns running, so does the Sequence, which will, when ticked again, either call the last running child or restart the Sequence, depending on its type: Sequence or ReactiveSequence.
Fallback
(a mirror of Sequence) Executes its children from left to right. If any child succeeds, the Fallback succeeds, skipping the rest. If a child returns failure, we tick the next child, unless it was the last one, in which case the Fallback returns failure. When a child returns running, the Fallback returns running, and, when ticked again, will tick either the last ticked child or the first one depending on its type: Fallback or ReactiveFallback.
Parallel
Runs/ticks multiple children simultaneously and succeeds or fails based on some success threshold (e.g., either all or some of the children must succeed for success).
Decorator nodes
perform a specific operation on a single child (e.g., the Invert node returns success when its child returns failure, and vice versa and returns running when the child returns running). Here is a non-exhaustive list of Decorator nodes: Repeat, ForceFailure, ForceSuccess, RetryUntilSuccessful, etc.
This is a very quick and shallow description of BT, and we invite the reader to check books (e.g., Colledanchise and Ögren (2018)) and online tutorials (e.g., https://www.behaviortree.dev/, which introduces the popular BehaviorTree.CPP), to get a complete picture of the BT “programming” ecosystem,
Note that the BT specifications are not “closed”. If most implementations propose the BT nodes described above, some offer additional Control nodes which implement more “specific” control algorithms (e.g., Nav2 Macenski et al. (2024) proposes Recovery, PipelineSequence, RoundRobin, …) or additional Decorator nodes (e.g., Nav2 proposes RateController, etc). Overall, BT are making explicit the control of the execution of the nodes. However, for the leaves of the tree (Action and Condition nodes), the specification remains minimal and silent on some features (e.g. can one pass arguments to nodes? returning values? asynchronous calls? time taken by the real execution? etc). Nevertheless, most implementations specify how these features are handled (e.g., with black board for variables, or in the C/C++ code called to implement them, multi-threading, etc).
Considering than more and more robotics applications, using BT, may be deployed in critical application (autonomous drones or vehicles, etc) or in environment with human (service robots), we need to be able to prove some safety properties on the BT, and to trust their execution will remain faithful to the programmer intentions. For this we believe we need to harness some formal models to the BT language and its execution engine to enable some formal offline and online validation and verification of BT.
2 State of the art and proposed approach
We split the state of the art in two parts, on one side, the approaches and the papers concerned with BT in robotic applications, on the other side, the ones which study formal models jointly with BT.
2.1 BT in robotics
BT are praised for their modularity, readability, scalability, flexibility and robustness, and supposedly being easy to debug and test. Even if these are questionable and somewhat subjective, one cannot deny their rising success and interest in robotics for autonomous navigation, human-robot interaction, manipulation tasks and multi-agent systems.
The book Colledanchise and Ögren (2018) covers most, if not all, aspects of BT in robotics. They make an extensive presentation of BT, how they compare to FSM, how they can be linked to the planning activity, etc. They mention the importance of safety and formalism, although not much is said on formal proof and verification.
In Iovino et al. (2022) the authors make a comprehensive and large survey of the topic of BT in AI and robotic applications. The existing literature is described and categorized based on methods, application areas and contributions, and the paper concludes with a list of open research challenges: explainable AI, human–robot interaction, safe AI, and the combination of learning and BT.
The work presented in Marzinotto et al. (2014) shows the equivalence between BT and Controlled Hybrid Dynamical Systems. Similarly the authors of Ögren and Sprague (2022) study how to deploy Behavior Trees in Robot Control Systems, and they propose an interesting formal analysis regarding convergence and regions of attraction.
The authors of Schulz-Rosengarten et al. (2024) address one of the BT “shortsight” and propose to add a cleaner communication extension, but lack formalism and proof on the LF. The input/output mechanism is inspiring.
As for evaluating BT, the authors of Gugliermo et al. (2024) propose a set of metrics (some static, some gathered from real runs), to evaluate some BT properties, as to evaluate and analyze them.
Implementation considerations
Deploying BT in robotic applications requires addressing implementation issues which may not be present in Video Game programming. In Colledanchise and Natale (2018) the authors present an original approach to handle parallelism and concurrency in BT (CBT) with execution progress and resources management. In Colledanchise and Natale (2021) , the same authors point out the issues on memory nodes (to avoid reevaluating), asynchronous action calls, parameters, halt (blocking or not), etc.
BT and planning
In many robotics architectures, BT is considered as the “acting” component of the decisional layer Ingrand and Ghallab (2017). This is the case for example in PlanSys2 Martín et al. (2021) (now part of the AIPlan4EU platform Micheli et al. (2025)) where the planner produces plans as BT which can be deployed for plan execution. It is also interesting to study how BT may also be extended to perform some planning. In Colledanchise et al. (2019) the authors propose a dynamic modification of BT for planning (planning with back chaining), so they can perform robust acting, without resorting to replanning. On another yet different type of planning/BT interaction, in Köckemann et al. (2023) planning is used to produce testing plans for BT, whose testing participates to increase the trust we put in these BT.
2.2 BT and formal models
This paper main subject is about BT and formal V&V, so we now examine the state of the art in this area.
In Klöckner (2013) the authors propose to interface BT mission plans and a simulation of the world using the description logic ($\mathcal{ALC(D)}$). So the description logic formal model acts as a safety check between the plan execution, and the simulation. Even if this does not provide a formal proof of the mission plans BT, it improves the trust we can have in these plans by formally validating their execution (in simulation) before deploying them in the real world.
The authors of Colledanchise et al. (2017) propose to synthesize correct by construction BT from an environment specification along the agent model and an objective expressed in LTL. From a standpoint, the approach is clearly sound and synthesizes correct BT, but requires the programmer to write LTL goal specifications to get started which may be seen as a deterrent to non formal “programmers”.
The last four approaches we present here have strong similarities with the one we propose.
In Biggar and Zamani (2020) the authors propose to synthesize LTL from BT and then show that the obtained model satisfies some LTL specifications. The paper goes in depth to explain the translation process, although it is not clear it can be automated, and a priori, the produced formal model cannot be directly executed in place of a regular BT engine.
The authors of Colledanchise et al. (2021) focus on the formalisation of the execution context of BT to be able to perform runtime verification. They propose channel systems to model the “surroundings” of the BT and then to check at runtime that some specifications, written in the SCOPE langage, are satisfied while the BT executes. So the approach, which is not limited to BT, provides a very strong and formal execution framework sitting between the BT and the robot, behave like a safety bag. The battery example they present has some similarity with the one we deploy on our UAV in section 6.1.
Similarly, the authors of Serbinowska et al. (2024) focus primarily on runtime verification of BT with contingency monitors (BTM) written with a DSL: BehaVerify. These monitors can be used to correct an undesirable behavior when it is detected and can handle LTL specifications. Yet, they can also check the BT at design time, by checking these BTM with model checking.
In Wang et al. (2024), the authors present an approach where they use the BIP formal framework to model BT and propose an implementation of their tool: xml2bip. They then use model checking (not D-Finder as the original BIP implementation did) to check for formal properties. Although some versions of BIP come with a runtime engine (e.g., the one used in Bensalem et al. (2011)), they do not yet propose a “fornal” execution of the BT with the BIP engine.
2.3 Proposed approach
Our approach has one main goal: to provide a formal semantics for BT, by translating it to a formal model, which can then be used offline to check formal properties, but also online to implement and enforce this semantics.
We propose to achieve this objective by following these steps.
  • Define a clear complete and unequivocal translation of all BT to a formal model in Fiacre Berthomieu et al. (2007, 2008).
  • The obtained BT formal model can be checked and analyzed to prove logical and temporal properties (LTL and CTL).
  • The same BT formal model can be linked to actual code and executed like other BT framework engines (e.g. Behaviortree.CPP, BT.py) do. This shows that the operational semantic of the BT formal model is the expected one, while guaranteeing the property proven offline.
Moreover, implementing this approach leads to some interesting side effects and features. It clarifies the BT semantics when needed, e.g., the wait/halt semantics when running nodes must be halted. It also enables time representation extensions and enriches the BT language with state variables and functions evaluation.
The rest of the paper is organized as follows. After introducing above the BT, the state of the art and our approach, we first present in Section 3 the Fiacre/Tina/Hippo formal suite we use. Section 4 introduces how each BT node is mapped in Fiacre (this is implemented in our BT2Fiacre tool [^1] The BT2Fiacre tool developed for this study can be downloaded from the repository: https://redmine.laas.fr/projects/bt2fiacre/pages/index.). Then Section  5 presents how these Fiacre BT nodes are put together to build the complete formal model of the whole BT, and we then show what are the type of formal properties one can prove offline but also at runtime. Two examples are presented: in Section 6.1, we introduce a drone controller written in BT for which we successfully deployed our approach, and; in Section 6.2, we show how the Nav2 BT Macenski et al. (2024) can be deployed with BT2Fiacre. A discussion in Section 7 reassesses the pros and cons of the BT2Fiacre tool and the use of Fiacre as an underlying formal language to provide a formal model and a formal semantics to BT, followed by future work section and the conclusion of the paper.
3 A Formal Framework for Offline and Runtime Verification: The Fiacre Language, Models, and Tools
While this paper does not aim to exhaustively present the formal framework we use, some terminology and explanations are essential for clarity and make the paper self contained. Readers interested in more details may consult the specific papers and websites cited below.  [^2] Note that a similar Fiacre presentation can be found in this paper Ingrand (2024a) (from the same author). We include it almost as is in this paper as to make the paper self contained, nevertheless, If the reviewers believe this section should be shorten and replace by a pointer to the other paper, this is perfectly fine for us.
3.1 Terminology, Models, Languages, and Tools
We define the following terms:
Time Petri Nets
Berthomieu and Diaz (1991) are an extension of traditional Petri nets where each transition has an associated time interval (typically $[0,\infty)$) specifying the time range within which a sensitized transition can be fired.
TTS
Time Transition Systems extend Time Petri nets by adding data-handling capabilities, allowing transitions to invoke data processing functions.
Tina
short for ”TIme Petri Net Analyzer,” is a toolkit for editing, simulating, and analyzing Petri nets, Time Petri nets, and TTS. Within this toolkit, sift and selt enable the construction of reachable state sets and the verification of LTL properties. [^3] https://projects.laas.fr/tina/index.php
Fiacre
stands for ”Intermediate Format for Embedded Distributed Component Architectures” (in French). It is a formally defined language designed to represent the behavioral and timing aspects of embedded and distributed systems for purposes of formal verification and simulation. Fiacre specifications can be compiled into a TTS using the frac compiler. [^4] https://projects.laas.fr/fiacre/index.php
H-Fiacre
extends the Fiacre language, making models “executable” by adding Event Ports and Tasks linked to C/C++ functions.
Hippo
is an engine for executing TTS resulting from the H-Fiacre specifications compilation (Hladik et al., 2021). [^5] https://projects.laas.fr/hippo/index.php
This framework has been applied across various projects and applications, [^6] https://projects.laas.fr/fiacre/papers.php including the validation and verification of functional components in our robotics experiments Dal Zilio et al. (2023), but also to the validation and verification of robotic skills programmed in ProSkill Ingrand (2024a).
3.2 Fiacre Semantics
[…]
Following these specifications, a component is defined by placing three process instances in parallel (line 1.56) and linking them through two ports (line 1.54). This example is simple by design, though the Fiacre language supports complex data types, bidirectional ports, local and global variables, conditions, switch/case statements, transition guards, and function calls (internal to Fiacre or external in C/C++ code) for advanced computation. More complex Fiacre specifications can be found in A and B.
Listing 1: Fiacre specification for a triple click detector (Fiacre offline version).
1process clicker [click:sync] is // synthesize clicks and sync them on its port at any time
2states wait_click, make_click
3
4from wait_click
5 wait [0, ...[; // wait any time from zero to infinity
6 to make_click
7
8from make_click
9 click; // issue a click sync on the Fiacre port
10 to wait_click
11
12process detect_triple_click [click:sync,triple_click:sync] is
13states wait_first, wait_second, wait_third, detected
14
15from wait_first
16 click; // first click
17 to wait_second
18
19from wait_second
20 select // we wait either
21 wait [0.2,0.2]; // exactly 0.2 second
22 to wait_first // then reset the detector
23 []
24 click; // or for the second click
25 to wait_third // whichever comes first
26 end
27
28from wait_third
29 select // again for the third click
30 wait [0.2,0.2];
31 to wait_first
32 []
33 click; // third
34 to detected
35 end
36
37from detected
38 triple_click; // sync on the triple_click port
39 to wait_first
40
41process triple_click_receiver[triple_click:sync] is
42states waiting_tc, received_tc
43
44from waiting_tc
45 triple_click; // just wait for a sync on this port
46 to received_tc
47
48from received_tc
49 /* do what needs to be done when a TC has been detected */
50 to waiting_tc
51
52component comp_tc is //we now specifiy the component
53
54port click:sync in [0,0], triple_click:sync in [0,0] // two ports
55
56par * in // 3 processes composed in parallel
57 detect_triple_click[click, triple_click] // process 1
58|| clicker[click] // process 2
59|| triple_click_receiver[triple_click] //process 3
60end
61
62comp_tc // this instantiates the component
63
64// some properties to check
65property ddlf is deadlockfree // deadlock free (TRUE)
66assert ddlf
67
68property cannot_receveice_tc is absent comp_tc/3/state received_tc
69assert cannot_receveice_tc // we cannot detect a triple click (FALSE)
Figure 1: The Fiacre processes modeling the Fiacre specification on Listing 1p1.
3.3 Offline Formal Verification
The frac compiler is used to compile the Fiacre specifications shown in Listing 1p1 into an equivalent TTS. With the sift tool from the Tina toolbox, one can then construct the system’s set of reachable states, and by using selt, the properties outlined in the initial Fiacre specifications, as well as any additional properties, can be verified. The Tina toolbox offers a variety of other tools that readers may explore for further analysis.
[…]
The verification approach used by the Tina tools relies on model checking, which can be affected by state explosion Clarke et al. (2012), potentially limiting its effectiveness. However, as demonstrated in section 6.1.1, the results from our example remain both insightful and non-trivial.
3.4 H-Fiacre Runtime Extensions
Although Fiacre was originally designed for offline verification, it has been extended with two primitives that enable runtime verification (Hladik et al., 2021). These extensions allow the model to connect with C/C++ functions that send events or execute commands, forming what we call H-Fiacre, to distinguish it from the base Fiacre language.
The purpose of the H-Fiacre runtime version is to make the model ”executable” in connection with real-world interactions.
Listing 2p2 (along with Figure 2p2) presents the executable version of the specification given in Listing 1p1.
Event Ports
are defined in the specification’s preamble (see line 2.1), linking a C function to a Fiacre port. In this case, the event click is linked to the c_click function in C/C++. When this port is one of the possible transitions (lines 1.161.24, and 1.33), the C/C++ function is called, and the port becomes active upon the function’s return. These C/C++ functions can accept and return values typed in Fiacre.
Tasks
are also defined in the preamble (see line 2.2), associating a task (in this case, report_triplec) with a C/C++ function (here c_report_triple_click), which is called asynchronously upon a start (see line 2.17). This enables the corresponding sync (line 2.21) once the C/C++ function completes. Values can be passed to the task at call time and returned when it completes.
Listing 2: H-Fiacre processes implementing a triple click detector.
1event click : sync is c_click // declare the Fiacre event port which transmits click
2task report_triplec () : nat is c_report_triple_click // The C/C++ function called by this task
3
4process detect_triple_click [triple_click:sync] is
5// this process is exactly the same than in the regular Fiacre version
6// only the click port is now an event port
7
8process triple_click_receiver[triple_click:sync] is
9states waiting_tc, received_tc, sync_report
10var ignore : nat
11
12from waiting_tc
13 triple_click;
14 to received_tc
15
16from received_tc // show an example of an external call
17 start report_triplec();
18 to sync_report
19
20from sync_report
21 sync report_triplec ignore; // wait until the call return
22 to waiting_tc
23
24component comp_tc is
25port triple_click:sync
26
27par * in
28 detect_triple_click[triple_click]
29|| triple_click_receiver[triple_click]
30end
31
32comp_tc
In this example, we replace the clicker process, which previously synchronized with click at any moment, with the click event port (highlighted in purple). Additionally, we introduce a task (report_triplec shown in light blue) to execute when synchronizing with a triple_click in the triple_click_receiver process. The remainder of the model remains unchanged, transforming it from a model specifying a triple-click detector to an actual program or controller that implements it. In this way, the specification itself becomes an executable program.
Figure 2: Illustration of the H-Fiacre program on Listing 2p2.
3.5 Runtime (Online) Verification
The H-Fiacre model, once compiled into TTS format using frac, is linked with the Hippo engine and the C/C++ functions needed to run it (e.g., c_click and c_report_triple_click). The Hippo engine executes the TTS model in real-time and initiates the appropriate C/C++ function calls (in separate threads) connected to event ports and tasks. Note that the properties checked offline also hold in the online version (as the reachable states set of the former include the one form the latter).
Additionally, the model can be extended with runtime verification properties by incorporating a monitoring process. For instance, if this controller is applied in a video game where rapid sequences of triple clicks are suspicious (indicating potential use of a cheating device), a new cheat_detector process could be added (shown on the left in Figure 2p2 and in Listing 3p3). This process could have three states and use a shared Boolean variable, cheat. This variable would be set to ‘false‘ upon transitioning to the waiting_second state in the detect_triple_click process, and switched to ‘true‘ upon reaching the detected state if the click sequence timing suggests non-human activity.
Listing 3: cheat_detector process detecting a cheating device by monitoring the cheat Boolean variable.
1process cheat_detector(&cheat:bool) is
2
3states state1, state2, cheat_detected
4
5from state1
6 on (not cheat); // guard on (not cheat)
7 to state2
8
9from state2 // cheat was set to false
10 select // either
11 wait [0.05,0.05]; // 50 ms elapsed
12 cheat := true; // reset the cheat variable
13 to state1 // go back to monitoring
14 []
15 on (cheat); // cheat became true again before the 50ms above.
16 to cheat_detected //caught cheating
17 end
18
19from cheat_detected
20 // the player is cheating, do what needs to be done.
21 to state1
[…]
This approach allows us to synthesize a controller that directly runs the specification. This dual capability is a major strength of the Fiacre framework: the same formal model can be verified offline and executed online. In practical terms, this means that the controller’s real-time behavior aligns with the initial model specifications, validating that the offline-verifiable properties are applied consistently in the live system. While observing expected behavior is a necessary, though not entirely sufficient, indicator of correctness, it significantly strengthens the link between specification and execution.
Moreover, if runtime behavior diverges from expectations, you can debug it as you would a typical program. From a formal perspective, the traces (or behavior patterns) of the H-Fiacre version (also called the Tina version) are contained within those of the Fiacre model (also called the Hippo version), ensuring consistency between the offline model and its runtime counterpart.
4 The mapping of BT in Fiacre
Before getting into the details of the produced formal models, we present on Figure 3p3 the overall workflow from BT to the formal executable version (top part in green), and the formal verifiable version and its analyzed properties report (bottom part in purple). One should keep in mind, that the BT programmers only provide the various BT in .btf format (like the one on Listing 4p4), the C/C++ codes which glue Action and Condition BT to the real robot commands and perception primitives (all in blue) and, optionally, LTL properties to verify, and monitors written in Fiacre (in slanted blue). The rest is fully synthesized and automatically compiled.
Figure 3: The BT2Fiacre (Hippo/Tina) workflow. Only the data in the blue boxes need to be provided by the programmer. The BT2Fiacre tool (in light red) synthesizes the two models, the rest is fully automated. In light green, the workflow for the Hippo runtime verification version, and in light purple, the workflow for the Tina offline verification version.
Listing 4: A simple drone survey BT. Note that for historical reasons and as we are reusing some of our existing tools, we do not use an XML syntax but rather a Lisp like syntax (called the .btf format), but this has no consequence on the content and the interpretation of BT (See Appendix C, Listing 10p10 and 11p11 for an example in both format).
1((BehaviorTree :name drone
2 (Sequence
3 (ParallelAll :wait 1 :halt 0 ; if wait is 1, will wait the running branch if one fails
4 (Action :ID start_drone)
5 (Action :ID start_camera))
6 (ReactiveSequence
7 (Fallback
8 (Condition :ID battery_ok) ; check if the battery is OK
9 (ForceFailure :ID fail ; if not, just land and fail
10 (Action :ID land)))
11 (Fallback
12 (Condition :ID localization_ok) ;same for localization
13 (ForceFailure :ID fail
14 (Action :ID land)))
15 (Sequence
16 (Action :ID takeoff :args (height 1.0 duration 0))
17 (Parallel :success 1 :wait 0 :halt 1 ; If the survey or the nav succeed, we are done.
18 (Action :ID camera_survey)
19 (Sequence
20 (Action :ID goto_waypoint :args (x -3 y -3 z 5))
21 (Action :ID goto_waypoint :args (x -1.5 y 3 z 5))
22 (Action :ID goto_waypoint :args (x 0 y -3 z 5))
23 (Action :ID goto_waypoint :args (x 1.5 y 3 z 5))
24 (Action :ID goto_waypoint :args (x 3 y -3 z 5))
25 (Action :ID goto_waypoint :args (x 3 y -3 z 5))))
26 (Action :ID goto_waypoint :args (x 0 y 0 z 5))
27 (Action :ID land)
28 (Action :ID shutdown_drone))))))
Figure 4: The graphical representation of the BT Listing 4p4.
As mentioned above, the semantics of BT is not formally defined. By translating BT to Fiacre, we define a formal semantics, hopefully consistent with the operational semantics people expect from BT.
Another goal we pursue is to model as much BT “additional information” as possible in the Fiacre model (for example if the variable used in BT can be modelled in Fiacre, the better).
We shall first present the overall mapping and then we will point out where the semantics had to be clarified and what are the “additional” information we gather in the Fiacre model.
4.1 The general mapping
Each BT node is automatically mapped in a Fiacre process (See Section 3.2), whose automata is modeled in accordance to its node type (control, decorator). Then, all these Fiacre processes are instantiated and composed in a component which provides a shared array of Fiacre records. This array BTnode[] has one element for each BT and is a shared variable between all the Fiacre processes. In the following figures BTnode[self] is the record for the current BT node. Each BTnode[] record has two fields of interest: caller and rstatus. caller is initialized at None and will be set to the BT node which ticks/calls it. It is set back to None when its execution (for the current tick) is finished. rstatus contains the last returned status (success, failure or running) of the BT node. rstatus may also be set to halt_me (by its caller), when a running BT node must be halted.
Figure 5: Preamble and postamble of BT nodes in Fiacre.
As shown on Figure 5p5, a BT node Fiacre process starts with a preamble (on Listing 8p8 or 9p9 from start_ to tick_node) with a guard on its activation on BTnode[self].caller (i.e. presumably its parent node has ticked its by setting its caller field). Follow a test to check if it has been asked to halt (in case it had previously returned running and now its parent asks it to halt). If it needs to run, it transitions to the tick_node state where the particular of this node type is handled.
Similarly, the postamble mostly consists of the four automata states (success, failure, running and halted), all returning the proper return status rstatus and then transitioning to a done state, in which the control is relinquished by setting caller to None (the parent node has a guard on this to check that the node has finished this tick).
In the following Action and Condition nodes in Fiacre are presented in their Tina version (i.e., the offline verification) but also Hippo version, (i.e. the runtime version). All other BT nodes Fiacre models are strictly the same between the two versions.
4.2 Condition Node
The Condition node is the simplest node, it only returns success or failure. The Tina version will just return either values (See Figure 6p6), while the Hippo version will call the external C/C++ function which performs the test and return its value (See Figure 7p7). The C/C++ function is expected to be fast and should not delay execution unnecessarily. Note that in our formalism, one can pass arguments to the C/C++ call. This is very much welcome to make the BT language more expressive and to some extends to have these values available in the formal model.
Figure 6: The Fiacre process modeling the Condition node for the verification (Tina) model.
Figure 7: The Fiacre process modeling the Condition node for the runtime (Hippo) model.
4.3 Action Node
The Action node is slightly more complex, as it can also return running. Indeed, actions may take some time and not return success or failure right away. To preserve the reactivity of the overall execution, and keep the tick short, it is often advised to return running while the action is still executing in its own thread. As a consequence, it means that an Action node may be halted (i.e. at some point, it is in a running state but its parent node wants to halt it). In the Tina version (See Figure 8p8), all the return values are possible, while in the Hippo version (See Figure 9p9), a Fiacre task is started calling the C/C++ action_task, and subsequent ticks will return running until the C/C++ action_task finishes and returns success or failure. An external C/C++ action_halt is also defined and is called if the action must be halted (it then returns failure). While action_task may take some time in its own thread, the action_halt is expected to be fast. In an Action node too, one may define and pass some arguments to the C/C++ function. Listing 8p8 in Appendix A show the Fiacre code for the takeoff action.
Figure 8: The Fiacre process modeling the Action node for the verification (Tina) model.
Figure 9: The Fiacre process modeling the Action node for the runtime (Hippo) model.
4.4 Sequence Nodes
The Sequence node ticks/executes its child in sequence as they succeed, until one fails. It then returns failure. If all succeed, it returns success. If one child returns running, the Sequence returns running. Upon return of the tick this particular child is ticked/called again.
Listing 9p9, shows the complete Fiacre code of a Sequence node with three children. The Fiacre process has a local variable next_seq initialized at 1, which holds which node will be ticked/called when the current node is ticked/called again. From the Tick Node state (See Figure 10p10), one proceeds to the $Child_{\texttt{next\_seq}}$ state, which ticks/calls the proper child. We wait until this child is done caller=None and check its returned status rstatus:
  • success, if it is the last child to success, we return success, otherwise, we proceed to the next child $Child_{\texttt{next\_seq+1}}$.
  • failure, we return failure.
  • running, we set next_seq to the proper value and return running.
There exist variants of Sequence: ReactiveSequence, SequenceWithMemory. They alter the way the sequence is ticked after failure or running are returned. For example ReactiveSequence always restarts the sequence from the beginning after one child has returned running. The goal here is not to list all the variants, just to give an hint on how these are transformed in Fiacre. Of course, the generated Fiacre code implements the proper semantics for each variant by properly setting the next_seq value.
Note that the Tina and the Hippo versions of the BT Fiacre process are exactly the same.
Figure 10: The Fiacre process modeling the Sequence node.
4.5 Fallback Nodes
The Fallback node executes/ticks its child in sequence as they fail, until one succeeds. It then returns success. If all fail, it returns failure. If one child returns running, it returns running. Upon return of the tick this child is again ticked/called. As shown on the Figure 11p11, it is the symmetric of the Sequence node, thus its behavior is just the mirror of the Sequence one described above.There is a variant of Fallback: ReactiveFallback. It alters the way the Fallback is restarted/reticked after running is returned. ReactiveFallback always restarts from the beginning after one child has returned running. This is achieved in Fiacre by properly setting the next_fb value.
Figure 11: The Fiacre process modeling the Fallback node.
4.6 Parallel Nodes
The Parallel node specifies how many children m out of all n children must succeed for the Parallel node to succeed (with ParallelAll they must all succeed: n = m). From this, we see that the implementation just ticks all the node, and then keep track of how many return failure, success or running (See Figure 12p12). If any final condition for success or failure is met (enough BTnode[child1].rstatus = success)888(enough… success) is true when success >= m or (enough BTnode[childn].rstatus = failure)999(enough… failure) is true when failure > n - m, then it proceeds to the corresponding state, otherwise, this node returns running. Note that Parallel nodes can lead to children being halted (if the condition for success or failure are met while some children are still running).
Figure 12: The Fiacre process modeling the Parallel node.
4.7 Decorator Nodes
Decorator nodes have only one child. When this child is done the transitions to success, failure or running depend on the particular decorator type (Inverter, ForceFailure, ForceSuccess, KeepRunningUntilFailure, RetryUntilSuccessful, RateController, Repeat, etc). For example, the Inverter (See Figure 13p13) one will swap the success and failure transitions, while leaving untouched the running one, Repeat will call its child a number of times, etc.
Figure 13: The Fiacre process modeling a Decorator (Inverter) node.
4.8 Semantics clarification and added Features
As mentioned earlier, the translation to Fiacre defines a formal semantics of BT. But there are particular choices which need to be clarified and specified. Many of these choices have already been identified Ghiorzi and Tacchella (2024), we try here to settle them with a formal proposition.
4.8.1 Halting running BT nodes
We have seen there are a number of situations where one may have to handle “orphan” running branches. They are not really orphans, but the situation is such that their parents have already decided the success or failure, independently of their own final outcome other than running. This happens when a Parallel fails (or succeeds) and there are still running children, similarly when a ReactiveSequence fails or a ReactiveFallback succeeds, one child may still be running. So for these nodes we added 2 keywords :wait and :halt to their specification. If :halt is true, we explicitly halt the still running children. Note that this is propagated all the way to leaves nodes, in particular Action nodes will explicitly halt whatever they are doing by calling their action_halt C/C++ function (see Section 4.3). If :wait is true, the node waits until all the children return something else than running. These choices can be discussed and modified, but at least, the semantics is clarified, defined and implemented in the formal model.
4.8.2 Ticks and BT root
[…]
The BT root is also responsible for keeping the BT alive. What should it do when its child return running, success or failure? Some implementation keep running on success, other do not. To clarify the model we produce, the root BT keeps running while its child returns running, and stops when it returns success or failure (e.g., in the drone experiment we present in Section 6.1, this is the expected behavior).
4.8.3 State Variables, Expression Evaluation and Node Status
BT tend to rely on “external” Actions and Conditions to compute values, tests them, etc. One of our objectives is to get as much as possible of the BT and its associated Condition and Action nodes modelled in Fiacre. The more we get, the more properties we can prove on the reachable state of the BT, and the more we control execution at runtime.
Hence we introduce state variables which can be used in the BT model and will end up in the Fiacre model as Fiacre variables.
For now, state variables can hold a natural number or an enumeration. See the example below with flight_level which can take an int between 0 and 3, or battery which can take three values (Good, Low and Critical) (note that for an enumeration, one can specify the acceptable transitions from one value to others).
We also added two new leaves nodes to the BT specification:
SetSV
nodes can be used to set a state variable passed with the :SV keyword, by calling the :ID function (defined in Fiacre). SetSV nodes only return success.
Eval
nodes evaluate the condition they hold and return success or failure.
Last, we also make available for evaluation the resulting status of any BT, e.g. (Eval (= camera_track.rstatus success)) will return success if the execution of the camera_track Action node was a success.
5 Deploying, Model Checking and Running the Fiacre BT
The final Fiacre model of the BT is produced by instantiating all the BT nodes Fiacre processes and combining them in parallel.
As shown on Figure 3p3 page 3 both Fiacre models (offline and online) are compiled in their respective TTS (i.e., a Time Petri Net with data) with the frac compiler.
5.1 Offline formal verification of BT
The Tina toolbox used here mostly rely on LTL (SE-LTL state/event LTL Chaki et al. (2004) to be more accurate). LTL already offers a rich language and a lot of flexibility to write and prove some logical and temporal properties on the formal BT. Using model checking, one can either check properties on the fly (reachability of a state), or show the absence of a state (but for this, the whole reachable state set has to be built). By default, the sift tool computes the reachable states set of the BT. The resulting .ktz file (binary kripke transition system) can then be analyzed with default or additional properties with the selt tools [^11] The Tina toolbox provides many tools, we just focus here on selt and sift..
For each BT node, we define and check some default properties: can the BT execute and complete, can it succeed, fail, or return running? Can it be halted? Most of these properties correspond to either a particular state in the fiacre model, or to some explicit value in the btnode[self].rstatus record. So the properties can be synthesized automatically as follow (by checking if their negation is reachable, i.e. we try to show that the state cannot be reached, and expect selt to return a counter example):
// For the Action_takeoff btnode
prove absent drone/12/state done
prove absent drone/12/value (btnode[takeoff].rstatus=success)
prove absent drone/12/value (btnode[takeoff].rstatus=failure)
prove absent drone/12/state halted
// the 12 is the index of the Action_takeoff btnode process instance in all the processes
// combined in parallel to build the drone component
One can also check more complex properties. For example that if we reach a particular state, then we will eventually reach another one.
We also modelled some examples found in some papers presented in the section 2.1.
In Biggar and Zamani (2020) the authors present an example of a Mars rover using unfolded solar panel to charge its battery, but should fold them when there is a storm. The same example is studied in  Wang et al. (2024) with the BT transformed to BIP on which they check the same properties. The equivalent BT in the .btf formalism can be found in Appendix D, Listing 12p12. After translation to Fiacre, we can build the set of reachable states:
sift $\neg$¬stats mars_rover.tts $\neg$¬rsd mars_rover.tts/mars_rover.ktz
20783 classe(s), 20275 marking(s), 72 domain(s), 103111 transition(s)
0.520s
and to prove the property (check that the robot cannot be hit by a storm while its solar panels are unfolded):
prove absent (mars_rover/3/state Unfolded and mars_rover/1/state Storm)
we get (as they do):
never (sv__panel__automata_1_sUnfolded /\ sv__meteo__automata_1_sStorm)
FALSE
selt finds a counter example, i.e., a state where this can happen (which is a problem and need to be fixed).
Similarly, in Wang et al. (2024) they develop an example with a train having a non nul speed while a door is open, we modelled it and reach the same formal proof results.
5.2 Online (or runtime) verification
The formal executable model is obtained by producing the TTS with the frac compiler, and then linking the result with the Hippo engine library and the C/C++ code implementing the Action and Condition Fiacre tasks and Fiacre externals (Section 4.3 and 4.2). In most case these C/C++ code use the client library which allows to call actions, services or check topics on ROS nodes Quigley et al. (2009); or to make requests to or read ports from G en oM modules Dal Zilio et al. (2023).
As mentioned earlier, our translation from .btf BT to Fiacre defines the formal semantics of the language. We have seen above that we can prove properties on this formal model. But executing the formal model with Hippo is also a way to “validate” that the formal semantics we defined is correct with respect to the operational semantics one expects from BT. In other words, the execution of the formal model produces what is expected by the original BT programmer, as if he/she were using a regular BT execution engine.
As seen on Figures 4p4 and 14p14, our BT show changing color while executing: white, the node has not yet been visited; dark blue, the tick is currently in this node; light blue, the tick has been passed to child(ren) node(s) below and the node is waiting for the children to return; yellow, the node has returned running; green, the node has returned success; red, the node has returned failure; purple, the node was running and has been halted by one of its parent nodes; pink, the node has been instructed to halt itself (and its running branches).
The white square node close to the root of the BT indicates the Hippo tick which is also the BT tick.
6 Some illustrating examples
We now illustrate our approach with two examples, a drone controller and the Nav2 navigation ROS2 stack.
6.1 An UAV surveying an area
We program an UAV to perform a survey mission with a BT. The functional layer of this experiment (Figure 14p14) has already been presented in Dal Zilio et al. (2023), but suffice to say that it provides robust localization, navigation, flight control and allows us to command the drone and its camera. It is deployed using the G en oM specification language (which also maps in the same formal framework to validate and verify the functional components Dal Zilio et al. (2023)).
Figure 14: Architecture of the drone experiment. On the left, all the functional components involved, on the right a Behavior Tree in charge of the overall mission.
Eight functional components are deployed for this experiment. The set of primitive commands available for the acting component are: start_drone, start_camera, shutdown_drone, takeoff, land, goto_waypoint, and camera_survey. Each of them has a corresponding Action node in the larger BT drone on listing 5p5 (e.g, takeoff (line 33), goto_waypoint (lines 39-45), etc),
The state variables handled by the model are battery: Good, Low and Critical (line 6); and flight_level: a natural number between 0 and 3 (line 1).
Listing 5: A more complex drone survey BT in the .btf format, with state variables declaration, SetSV and Eval nodes. The :=assign:=:= is the assignment operator, and $\sim$ is the logical negation (See also Figure 15p15).
1((defsv fls ; the flight level of the drone
2 :init 0
3 :min 0
4 :max 3)
5
6(defsv battery ; the battery level of the drone
7 :states (Good Low Critical) ;Self explanatory
8 :init Good
9 :transitions :all)
10
11(BehaviorTree :name drone
12 (Sequence
13 (ParallelAll :wait 1 :halt 0
14 (Action :ID start_drone)
15 (Action :ID start_camera))
16 (ReactiveSequence :halt 1
17 (Sequence
18 (Fallback
19 (ForceFailure :ID fail
20 (SetSV :ID measure_battery :sv battery))
21 (Eval (~ (= battery critical)))
22 (ForceFailure :ID fail_mission
23 (Action :ID land)))
24 (Eval (~ (= battery critical)))
25 (Fallback
26 (Eval (= battery good))
27 (Action :ID goto_waypoint :args (x $\neg$1 y $\neg$1 z 0.5)))) ; goto charging station
28 (Fallback
29 (Condition :ID localization_ok)
30 (ForceFailure :ID fail
31 (Action :ID land)))
32 (Sequence
33 (Action :ID takeoff :args (height 1.0 duration 0))
34 (Parallel :success 1 :wait 0 :halt 1 ; if the tracking or the nav success... we are done.
35 (Action :ID camera_tracking :name camera_track)
36 (Repeat :repeat 3
37 (Sequence
38 (Eval (:= fls (+ 1 fls)))
39 (Action :ID goto_waypoint :args (x $\neg$3 y $\neg$3 z (* 2 \$fls)))
40 (Action :ID goto_waypoint :args (x $\neg$1.5 y 3 z (* 2 \$fls)))
41 (Action :ID goto_waypoint :args (x 0 y $\neg$3 z (* 2 \$fls)))
42 (Action :ID goto_waypoint :args (x 1.5 y 3 z (* 2 \$fls)))
43 (Action :ID goto_waypoint :args (x 3 y $\neg$3 z (* 2 \$fls)))
44 (Action :ID goto_waypoint :args (x 3 y $\neg$3 z (* 2 \$fls))))))
45 (Action :ID goto_waypoint :args (x 0 y 0 z 5))
46 (Action :ID land)
47 (Action :ID shutdown_drone)
48 (Eval (= camera_track.rstatus success)))))))
Figure 15: The graphical representation of the BT Listing 5p5.
Calling BT2Fiacre results in a Fiacre model with two processes for the state variables, and thirty eight processes for the BT nodes. [^12] The resulting code can be found in the examples subdirectory of https://redmine.laas.fr/projects/bt2fiacre/repository.
6.1.1 Offline verification results
To compute the reachable states set of the drone BT (Listing 5p5) we call the sift tool:
sift $\neg$¬stats drone.tts drone.tts/drone.ktz
49 196 302 classe(s), 49 145 735 marking(s), 152 domain(s), 267 702 880 transition(s)
4552.704s
which takes around 1 hour and 16 minutes to build on an Intel(R) Xeon(R) Silver 4110 CPU @ 2.10 GHz with 256 GB of memory. The resulting .ktz has $49196302$ classes, $49145735$ markings, $152$ domains, $267702880$ transitions.
All the default properties presented in Section 5.1 have been checked and all the results are the expected ones. Most of them take between less than a second and 300 seconds to be verified. As expected, the ones for which the model checker finds a counter example are usually quick to be verified while the ones which require to explore the whole set take more time.
Among the default proven properties, somes are still interesting to comment. For example, despite the ParallelAll on the start_drone and the start_camera, they cannot be halted (because the ParallelAll has :halt 0). Unlike the BT children of the ReactiveSequence which has :halt 1, they can all be halted (except the first branch) [^13] As only a branch returning running on the previous tick may be halted if one branch on its left fails, hence, never the first branch..
But on the top of the default properties, we can add new ones specific to this particular experiment. Can we prove that if the battery is Critical then the drone will attempt to land, and if it is Low it will return to the charging station? Can we show that if the camera survey fails then the whole drone mission will fail? Can the drone fly higher than 6 meters?
These properties can easily be written in Fiacre pointing to states of BT Fiacre nodes (e.g., tick_node, failure), values of state variable (e.g. fls, battery), or even returned status from executed BT (e.g., btnode[localization_ok_btn18].rstatus = failure) [^14] The ,$\Box$,$\Rightarrow$,$\diamondsuit$ , ⇒ , ♢ operators have the usual LTL semantics.:
property attempt_to_land_if_battery_Critical is
ltl ($\Box$ (drone/5/value (battery=Critical)) $\Rightarrow$
$\diamondsuit$ (drone/8/state tick_node)) // land_btn12
property attempt_to_go_to_charging_station_if_battery_Low is
ltl ($\Box$ (drone/12/value (battery=Low)) $\Rightarrow$
$\diamondsuit$ (drone/13/state tick_node)) // goto_waypoint_btn16
property attempt_to_land_if_localization_failure is
ltl ($\Box$ (drone/16/value (btnode[localization_ok_btn18].rstatus = failure)) $\Rightarrow$
$\diamondsuit$ (drone/17/state tick_node)) // land_btn20
property camera_fails_implies_mission_fails is
ltl ($\Box$ (drone/21/state failure) $\Rightarrow$ // camera_tracking_camera_track
$\diamondsuit$ (drone/39/state failure)) // BehaviorTree1_drone
property fly_not_higher_than_6m is absent (drone/35/value (fls > 3))
The result are proven true for all these properties:
operator attempt_to_land_if_battery_Critical : prop
TRUE
0.001s
operator attempt_to_go_to_charging_station_if_battery_Low : prop
TRUE
0.001s
operator attempt_to_land_if_localization_failure : prop
TRUE
0.001s
operator camera_fails_implies_mission_fails : prop
TRUE
0.001s
operator fly_not_higher_than_6m : prop
TRUE
242.462s
More interestingly, we can also take advantage of Fiacre/Tina timed models. For example, if one considers the one tick per node semantics (See Section 4.8.2), we can prove that the a critical battery will always leads to a landing within a $[0,2]$ ticks interval (everything else considered in the BT):
property attempt_to_land_if_battery_Critical_timed is
(drone/5/value (battery=Critical)) leadsto (drone/8/state tick_node) within [0,2]
which results in:
operator attempt_to_land_if_battery_Critical_timed : prop
TRUE
49.170s
6.1.2 Runtime verification results
The Drone BT execution by Hippo runs as expected, the survey mission is executed, and by adding random fault (on battery level, or failing the localization_ok Condition node), we show that the drone behaves as expected (perform land to prevent further problems). If the survey finds the object, it returns success and the navigation is halted (because the Parallel is :success 1 :wait 0 :halt 1. Note that at the end, we check with an Eval mode, that the returned status of the camera_track Action node is success. So when the overall Drone BT completes, it will return success if and only if the target was found, failure otherwise.
Again, the fact that the BT formal model execution exhibit what is expected from the original programmer is a good sign that the formal semantics is consistent with the operational one, and that the offline proofs we did on the very same model hold for the original BT. [^15] Some videos of these runs and the BT animation are available here: https://redmine.laas.fr/projects/bt2fiacre/pages/index#yet-another-more-complex-example-drone3.
From a performance point of view, the overhead of the Hippo execution is negligible. After all, the Hippo engine is a Petri Net (TTS) “execution” engine, and has 497 places and 1113 transitions and manages up to 7 task execution threads (one for each Action BT node).
Figure 16: The screen dump of the drone BT mission 5p5 executing in Hippo. This shows that the top BehaviorTree and its child Sequence child have passed the tick to the ReactiveSequence currently executing while its Sequence has returned running from its Land child, while the camera_track has returned success, as a result, the Parallel has halted the Repeat and its children. Note the tick counter near the root which indicates the mission is in its 20th second of execution.
6.2 ROS2 Nav2 BT
One of the most popular navigation stacks in robotics is the Nav2 one. It is now part of ROS2 and is proposed with a number of BT to implement it Macenski et al. (2024). So to also properly test our approach we implemented the Nav2 BT in our framework. Note that these BT define some new Decorator and Control nodes: Recovery, PipelineSequence, RateController, RoundRobin, etc. They were implemented in our BT2Fiacre tool and are being translated to Fiacre along the regular ones. One interesting aspect of our approach is that the operational semantics of these new nodes can be proven in our formal semantic of Fiacre using the Tina toolbox. For example, one can formally prove that our implementation of Recovery is correct by showing that the second node can only be called when the first one has returned failure, and that the first one can only be called again when the second has returned success, etc.
We are able to check the reachable state of the Nav2 BT presented in Appendix, Listing 11p11 and Figure 18p18.
sift $\neg$¬stats ros2$\neg$¬nav2.tts ros2$\neg$¬nav2.tts/ros2$\neg$¬nav2.ktz
171 656 098 classe(s), 171 656 098 marking(s), 66 domain(s), 422 970 818 transition(s)
5776.872s
selt ros2$\neg$¬nav2.tts/ros2$\neg$¬nav2.ktz ros2$\neg$¬nav2.tts/ros2$\neg$¬nav2.ltl $\neg$¬b
Selt version 3.8.0 $\neg$$\neg$ 05/02/24 $\neg$$\neg$ LAAS/CNRS
ktz loaded, 171 656 098 states, 422 970 818 transitions
2212.781s
[…]
More interestingly is to check that our implementation of the added BT (e.g., Recovery, RoundRobin, etc) is correct.
Recovery
The Recovery node [^16] https://docs.nav2.org/behavior_trees/overview/nav2_specific_nodes.html is a control flow node with two children. It returns success if and only if the first child returns success. The second child will be executed only if the first child returns failure. If the second child succeeds, then the first child will be executed again. The user can specify how many :retry times the recovery actions should be taken before returning failure.
We define a simple generic example:
Listing 6: A simple Recovery node.
1((BehaviorTree :name bt_recovery
2 (Recovery :num_retries 1 :name recovery
3 (Action :name action)
4 (Action :name recov))))
for which the model checker selt is able to prove the following properties:
// Action_action // 1
// Action_recov // 2
// Recovery_recovery // 3
// BehaviorTree_bt_recovery // 4
property failure_recov_leads_to_failure is // TRUE if the recovery action fails, than the recovery fails
(bt_recovery/2/state failure) leadsto (bt_recovery/3/state failure) within [0,0]
property failure_action_leads_to_recovery_failure is // FALSE because retry is still 0 (not retried yet)
(bt_recovery/1/state failure and bt_recovery/3/value (retry = 0)) leadsto (bt_recovery/3/state failure) within [0,0]
property failure_action_leads_to_recovery_failure is // TRUE because we already retied
(bt_recovery/1/state failure and bt_recovery/3/value (retry = 1)) leadsto (bt_recovery/3/state failure) within [0,0]
property success_action_leads_to_success is // TRUE
(bt_recovery/1/state success) leadsto (bt_recovery/3/state success) within [0,0]
We present here a proof on a simple instance, but similar proofs can be made on an instance of Recovery embedded in a larger BT, for example the ones in the Nav2 BT in Appendix, Listing 11p11 and Figure 18p18.
RoundRobin
The RoundRobin node [^17] https://docs.nav2.org/behavior_trees/overview/nav2_specific_nodes.html is a control node which ticks its children in a round robin fashion until a child returns success, in which case the parent node will also return success. If all children return failure so will the parent RoundRobin.
Similarly, we define a simple generic example:
Listing 7: A simple RoundRobin node.
1((BehaviorTree :name bt_roundrobin
2 (KeepRunningUntilFailure :name kr
3 (RoundRobin :name RR
4 (Action :name A1)
5 (Action :name A2)
6 (Action :name A3)
7 (Action :name A4)))))
for which the model checker is able to prove the following properties:
// Action_A1 // 1
// Action_A2 // 2
// Action_A3 // 3
// Action_A4 // 4
// RoundRobin_RR // 5
property failure_a1_leads_to_failure is // FALSE a failure of one child does not necessary leads to rr failure
(bt_roundrobin/1/state failure) leadsto (bt_roundrobin/5/state failure) within [0,1]
property failure_a4_leads_to_a1_ticked is // TRUE a failure of a child with less than 3 failure so far leads to the next child to be ticked
((bt_roundrobin/4/state failure ) and (bt_roundrobin/5/value (failed < 3))) leadsto (bt_roundrobin/1/state tick_node) within [0,0]
property success_a2_leads_a3_ticked_exp_true is // TRUE success of a child leads to the next one to be ticked upon rr reticked
(bt_roundrobin/2/state success) leadsto (bt_roundrobin/3/state tick_node) within [0,3] // 3 because the tick has to go all the way up
property success_a2_leads_rr_success_exp_true is // TRUE
(bt_roundrobin/2/state success) leadsto (bt_roundrobin/5/state success) within [0,0]
Similar proofs can be made with any BT nodes to show that our Fiacre implementation satisfies the expected formal properties defining the BT node considered.
6.2.1 Runtime
All the Nav2 BT have been translated to .btf and run with Hippo and simulated ROS2 actions. One of the features we added and we believe goes beyond the “nice graphical touch” aspect, is to dynamically color the BT nodes while they execute. Of course, there are text traces of the Hippo engine running the BT, but being able to follow the execution ticks as well as the last returned status for each BT is rather informative. Some videos of these runs and the BT animation are available here: https://redmine.laas.fr/projects/bt2fiacre/pages/index#playing-with-ros2-nav2-bt. Figure 17p17 shows the Nav2 BT (See Listing 11p11 and Figure 18p18) executed by Hippo.
Figure 17: The screen dump of the Nav2 BT (See Listing 11p11 and Figure 18p18) executing.
7 Limits, discussion, future work, and conclusion
Before concluding, we propose to examine the limits of our work, to discuss its pros and cons and consider future work.
7.1 Limits
Some limits of our work are mostly due to choices currently made, which could be reconsidered, if needed. For example, we do not implement a black board to handle variables. Instead we propose to handle BT variables in Fiacre directly. This has the advantage of being able to include these variables in the formal model, hence in the proof and the runtime verification (e.g. the fls flight level in the drone experiment). The disadvantage is that only Fiacre supported types can be used.
An area where our approach would suffer, is when BT are dynamically modified, or transformed. For example in Section 2.1 we consider BT used for planning. In these approaches, BT are being modified on the fly, clearly, we could not perform model checking on the fly, yet assuming the dynamic of the planner remains slow (below a second) we could still synthesize the model, compile it and jump in its execution on the fly.
Although all our examples are based on one BT, absolutely nothing prevent us from having multiple BT executing together. yet from a verification point of view, this would probably lead to intractable model as models running in parallel tend to multiply the size of their reachable states set, to account for all the possible transition interleavings.
This brings us to the main limiting factor of our approach. Offline verification with model checking may lead to large intractable reachable states set. This is a well known limit of these approaches and we do not have any magic bullet. Yet we have seen that we can verify reasonably complex robotics BT skills, that we can prove “individual” BT node behavior (e.g. the ones added in Nav2), and that we can also handle complex properties on missions written with BT.
7.2 Discussion
Our approach to improve the trust one can put in BT completely relies on a on a well established formal language and framework (Fiacre and TTS), a formal verification toolbox (Tina to check LTL, CTL, properties) and Hippo, a runtime engine able to execute the formal model. We have seen in Section 2.2 that there are other approaches which transform BT in a formal framework, or harness a formal model around BT, but to our knowledge, none of them support the automatic translation of BT to an equivalent formal model and the execution the same formal model in place of the original BT. This is a very strong argument as the same formal model, which clarifies and specifies the formal semantics of BT, can be used to prove properties and also execute and show while running that it properly implements the expected operational semantic, and knowing that the proof made offline also holds online. As for taking into account the execution environment, it is taken into account in the perception provided by Condition and Action nodes.
Moreover, we have shown that the formal model of the various BT nodes (even the ones added for new experiments) can also be formally checked with our toolbox. Moreover, even if Fiacre and Tina do not provide per se a timed logic (e.g. TCTL or MITL), the toolbox provides some patterns Abid et al. (2014) which allow the verification of properties with explicit time (e.g. to prove that at most x units of time will elapse between these two states).
Moreover, Fiacre offers some features which could be valuable to be added to BT. Explicit time representation is the foremost feature which could come handy if added to the BT . Indeed, robot planning, acting and control is usually handling explicit time representation. Control loops have frequency, plans have duration and deadline, actions take some time to execute, etc. So it would be perfectly logical to add explicit time representation to BT. As a consequence, this would also have an impact on the tick semantic which would probably become more an execution token than a time elapsing counter.
Another domain where our approach and Fiacre could offer some valuable addition is with environment modelling. BT are intrinsically embedded operational models, they really exists to be executed in the real world (or in simulation). But now that we can consider model checking them, one needs to take into account (when possible) the outcomes of actions or conditions in the environment. Without any particular information, the model checker consider all possibilities (success, failure or running), but the real world often prove to be more “constrained”, and one could consider a more accurate and explicit model of the environment (e.g., with external asynchronous events, state variable values changes, etc).
Parallelism is clearly allowed and properly modelled in our framework. Moreover the model checker can verify properties even if it needs to consider all the possible execution interleavings of parallel nodes (but at some cost). Meanwhile, the semantic of transition in Fiacre is such that resource checking and reservation on a transition are atomic. As a consequence, resources management, even considering parallelism comes for free with Fiacre.
Last, our approach has been implemented in a software suite Ingrand (2024b) which has been applied to many BT examples (available here: https://redmine.laas.fr/projects/bt2fiacre/pages/index)
7.3 Future work
Considering the tight links between the Fiacre framework and our approach/tool BT2Fiacre, it is clear that any improvement in the former may also improve the latter. The Fiacre developers are considering extending the data types handled in the language by adding rational numbers and strings. Both types would still allow model checking and enable the deployment of richer state variables in our .btf format.
Similarly, there are many existing features in Fiacre which could be used in .btf BT (some have already been used in ProSkill Ingrand (2024a), another robotics acting language mapping to Fiacre). As already mentioned in Section 7.2 above, time would be a valuable addition to model, with time interval $[min,max]$, how long an Action is expected to take, or to wait a given amount of time before returning from a Wait node, etc. Not only would this enrich the .btf format and language, but it would also be taken into account by the Tina model checking tools, and enforced by the Hippo engine.
External asynchronous events and state variables asynchronous value changes are features also available in Fiacre and its tools. This again could enrich the .btf format and allow the programmer to account for “uncontrollable” state transitions.
7.4 Conclusion
BT are more and more popular in robotics. To encourage their deployment and improve the trust one has in robotic applications using BT, we propose an approach and an automatic tool to transform any BT in a formal model with a formal semantics. The resulting models can then be used offline with model checking to prove some properties of the BT, but also linked to the real actions and perceptions of the robots and execute online on the robot. Of course, this can be deployed in BT applications outside of robotics, and also participates to define extensions to BT and to better formalize them.
Acknowledgement
We thank Bernard Berthomieu, Silvano Dal Zilio and Pierre-Emmanuel Hladik for their help while developing and deploying the work presented here.
Appendix A A BT Action H-Fiacre process
Listing 8: The H-Fiacre process specification of the takeoff BT Action node.
1process btnode_takeoff_btn21 (&btnode: btnode_array, &fls: sv_fls, &battery: sv_battery) is
2
3states start_, tick_node, success, failure, halt, halted, running, error,
4 Action_takeoff, dispatch, Action_takeoff_sync, done
5
6var callb: bool,
7 report_halted:bool := false, ret_val: ret_status
8
9from start_
10 wait [0,0];
11 on (btnode[takeoff_btn21].caller $\diamondsuit$ None); // Wait until we are called
12 report_halted := false;
13 if (btnode[takeoff_btn21].rstatus = halt_me) then // are we being instructed to halt?
14 report_halted := true;
15 to halt
16 end;
17 // Btnode Action ’takeoff_btn21’ has been called (:height 1.000000 :duration 0 )
18 btnode[takeoff_btn21].rstatus := no_ret_status; // just initializing the rstatus
19to tick_node
20
21from halt
22 wait [0,0];
23 // Btnode Action ’takeoff_btn21’ is being halted (:height 1.000000 :duration 0 )
24 callb := Fiacre_Action_takeoff_halt (btnode[takeoff_btn21]); //This call the external which
25to halted // halts the action
26
27from tick_node
28 wait [0,0];
29 to Action_takeoff
30
31from Action_takeoff
32 // synthesized action arg index 2
33 btnode[takeoff_btn21].ArgIndex := 2;
34 // Btnode Action ’takeoff_btn21’ calling its action (start task)
35 start Fiacre_Action_takeoff_task (btnode[takeoff_btn21]); // this call the Fiacre task
36to Action_takeoff_sync // which handles this action.
37
38from Action_takeoff_sync
39 sync Fiacre_Action_takeoff_task ret_val; // wait until the Fiacre task returns
40 // Btnode Action ’takeoff_btn21’ returned (sync task)
41 to dispatch
42
43from dispatch
44 wait [0,0]; // we dispatch to the proper state according to the return values
45 if (ret_val = success) then
46 to success
47 elsif (ret_val = failure) then
48 to failure
49 elsif (ret_val = running) then
50 to running
51 else
52 to error // a priori unreachable
53 end
54
55from success
56 wait [0,0];
57 // Btnode Action ’takeoff_btn21’ (:height 1.000000 :duration 0 ) returns success.
58 btnode[takeoff_btn21].rstatus := success;
59 to done
60
61from failure
62 wait [0,0];
63 if (report_halted) then // this is mostly for traces
64 // Btnode Action ’takeoff_btn21’ (:height 1.000000 :duration 0 ) returns halted failure.
65 null
66 else
67 // Btnode Action ’takeoff_btn21’ (:height 1.000000 :duration 0 ) returns failure.
68 null
69 end;
70 btnode[takeoff_btn21].rstatus := failure;
71 to done
72
73from halted
74 wait [0,0];
75 // Btnode Action ’takeoff_btn21’ (:height 1.000000 :duration 0 ) has been halted.
76 report_halted := true;
77 to failure
78
79from running
80 wait [0,0];
81 // Btnode Action ’takeoff_btn21’ (:height 1.000000 :duration 0 ) returns running.
82 btnode[takeoff_btn21].rstatus := running;
83 to done
84
85from done
86 wait [0,0];
87 // Btnode Action ’takeoff_btn21’ returning control to caller and back to ’start_’
88 btnode[takeoff_btn21].caller := None; // we relinquish the tick and go back waiting
89to start_
Appendix B A BT Sequence H-Fiacre process
Listing 9: The H-Fiacre process specification of a BT Sequence node.
1process btnode_Sequence13_btn6 (&btnode: btnode_array) is
2
3states start_, tick_node, success, failure, halt, halted, halt_wait, running, error,
4 Fallback15_btn7, Fallback15_btn7_done, Eval25_btn12, Eval25_btn12_done,
5 Fallback27_btn13, Fallback27_btn13_done, done
6
7var next_seq: 1..3 := 1
8
9from start_
10 wait [0,0];
11 on (btnode[Sequence13_btn6].caller $\diamondsuit$ None);
12 if (btnode[Sequence13_btn6].rstatus = halt_me) then to halt end;
13 // Btnode Sequence ’Sequence13_btn6’ has been called
14 btnode[Sequence13_btn6].rstatus := no_ret_status;
15to tick_node
16
17from halt
18 wait [0,0];
19 // Btnode Sequence ’Sequence13_btn6’ is being halted
20 if (btnode[Fallback15_btn7].rstatus = running) then
21 // Btnode Sequence ’Sequence13_btn6’ halting Fallback ’Fallback15_btn7’
22 btnode[Fallback15_btn7].rstatus := halt_me;
23 btnode[Fallback15_btn7].caller := caller_Sequence13_btn6;
24 to halt_wait
25 end;
26 if (btnode[Eval25_btn12].rstatus = running) then
27 // Btnode Sequence ’Sequence13_btn6’ halting Eval ’Eval25_btn12’
28 btnode[Eval25_btn12].rstatus := halt_me;
29 btnode[Eval25_btn12].caller := caller_Sequence13_btn6;
30 to halt_wait
31 end;
32 if (btnode[Fallback27_btn13].rstatus = running) then
33 // Btnode Sequence ’Sequence13_btn6’ halting Fallback ’Fallback27_btn13’
34 btnode[Fallback27_btn13].rstatus := halt_me;
35 btnode[Fallback27_btn13].caller := caller_Sequence13_btn6;
36 to halt_wait
37 end;
38 to halted
39
40from halt_wait
41 on ((btnode[Fallback15_btn7].caller = None) and
42 (btnode[Eval25_btn12].caller = None) and
43 (btnode[Fallback27_btn13].caller = None));
44 to halted
45
46from tick_node
47 wait [0,0];
48 if (next_seq = 1) then to Fallback15_btn7 end;
49 if (next_seq = 2) then to Eval25_btn12 end;
50 if (next_seq = 3) then to Fallback27_btn13 end;
51 to error
52
53from Fallback15_btn7
54 wait [0,0];
55 // Btnode Sequence ’Sequence13_btn6’ calling Fallback ’Fallback15_btn7’
56 btnode[Fallback15_btn7].caller := caller_Sequence13_btn6;
57 to Fallback15_btn7_done
58
59from Fallback15_btn7_done
60 wait [0,0];
61 on (btnode[Fallback15_btn7].caller = None);
62 // Btnode Sequence ’Sequence13_btn6’ getting control back from Fallback ’Fallback15_btn7’
63 if (btnode[Fallback15_btn7].rstatus = success) then
64 to Eval25_btn12
65 elsif (btnode[Fallback15_btn7].rstatus = failure) then
66 next_seq := 1;
67 to failure
68 elsif (btnode[Fallback15_btn7].rstatus = running) then
69 next_seq := 1;
70 to running
71 else
72 to error
73 end
74
75from Eval25_btn12
76 wait [0,0];
77 // Btnode Sequence ’Sequence13_btn6’ calling Eval ’Eval25_btn12’
78 btnode[Eval25_btn12].caller := caller_Sequence13_btn6;
79 to Eval25_btn12_done
80
81from Eval25_btn12_done
82 wait [0,0];
83 on (btnode[Eval25_btn12].caller = None);
84 // Btnode Sequence ’Sequence13_btn6’ getting control back from Eval ’Eval25_btn12’
85 if (btnode[Eval25_btn12].rstatus = success) then
86 to Fallback27_btn13
87 elsif (btnode[Eval25_btn12].rstatus = failure) then
88 next_seq := 1;
89 to failure
90 elsif (btnode[Eval25_btn12].rstatus = running) then
91 next_seq := 2;
92 to running
93 else
94 to error
95 end
96
97from Fallback27_btn13
98 wait [0,0];
99 // Btnode Sequence ’Sequence13_btn6’ calling Fallback ’Fallback27_btn13’
100 btnode[Fallback27_btn13].caller := caller_Sequence13_btn6;
101 to Fallback27_btn13_done
102
103from Fallback27_btn13_done
104 wait [0,0];
105 on (btnode[Fallback27_btn13].caller = None);
106 // Btnode Sequence ’Sequence13_btn6’ getting control back from Fallback ’Fallback27_btn13’
107 if (btnode[Fallback27_btn13].rstatus = success) then
108 next_seq := 1;
109 to success
110 elsif (btnode[Fallback27_btn13].rstatus = failure) then
111 next_seq := 1;
112 to failure
113 elsif (btnode[Fallback27_btn13].rstatus = running) then
114 next_seq := 3;
115 to running
116 else
117 to error
118 end
119
120from success
121 wait [0,0];
122 // Btnode Sequence ’Sequence13_btn6’ returns success.
123 btnode[Sequence13_btn6].rstatus := success;
124 to done
125
126from failure
127 wait [0,0];
128 // Btnode Sequence ’Sequence13_btn6’ returns failure.
129 btnode[Sequence13_btn6].rstatus := failure;
130 to done
131
132from halted
133 wait [0,0];
134 // Btnode Sequence ’Sequence13_btn6’ has been halted.
135 btnode[Sequence13_btn6].rstatus := failure;
136 to done
137
138from running
139 wait [0,0];
140 // Btnode Sequence ’Sequence13_btn6’ returns running.
141 btnode[Sequence13_btn6].rstatus := running;
142 to done
143
144from done
145 wait [0,0];
146 // Btnode Sequence ’Sequence13_btn6’ returning control to caller and back to ’_start’
147 btnode[Sequence13_btn6].caller := None;
148to start_
Appendix C A ROS2 Nav2 BT and its BTF equivalent.
Listing 10: One of the ROS2 Nav2 Behavior Tree (Navigate To Pose With Replanning and Recovery) Macenski et al. (2024).
1<root main_tree_to_execute="MainTree">
2 <BehaviorTree ID="MainTree">
3 <RecoveryNode number_of_retries="6" name="NavigateRecovery">
4 <PipelineSequence name="NavigateWithReplanning">
5 <RateController hz="1.0">
6 <RecoveryNode number_of_retries="1" name="ComputePathToPose">
7 <ComputePathToPose goal="{goal}" path="{path}" planner_id="GridBased"/>
8 <ReactiveFallback name="ComputePathToPoseRecoveryFallback">
9 <GoalUpdated/>
10 <ClearEntireCostmap name="ClearGlobalCostmap¬$\neg$¬Context"
11 service_name="global_costmap/clear_entirely_global_costmap"/>
12 </ReactiveFallback>
13 </RecoveryNode>
14 </RateController><
15 <RecoveryNode number_of_retries="1" name="FollowPath">
16 <FollowPath path="{path}" controller_id="FollowPath"/>
17 <ReactiveFallback name="FollowPathRecoveryFallback">
18 <GoalUpdated/>
19 <ClearEntireCostmap name="ClearLocalCostmap¬$\neg$¬Context"
20 service_name="local_costmap/clear_entirely_local_costmap"/>
21 </ReactiveFallback>
22 </RecoveryNode>
23 </PipelineSequence>
24 <ReactiveFallback name="RecoveryFallback">
25 <GoalUpdated/>
26 <RoundRobin name="RecoveryActions">
27 <Sequence name="ClearingActions">
28 <ClearEntireCostmap name="ClearLocalCostmap¬$\neg$¬Subtree"
29 service_name="local_costmap/clear_entirely_local_costmap"/>
30 <ClearEntireCostmap name="ClearGlobalCostmap¬$\neg$¬Subtree"
31 service_name="global_costmap/clear_entirely_global_costmap"/>
32 </Sequence>
33 <Spin spin_dist="1.57"/>
34 <Wait wait_duration="5"/>
35 <BackUp backup_dist="0.15" backup_speed="0.025"/>
36 </RoundRobin>
37 </ReactiveFallback>
38 </RecoveryNode>
39 </BehaviorTree>
40</root>>
Listing 11: The .btf version of the ROS2 Nav2 BT above (Listing 10p10).
1((BehaviorTree :ID MainTree ; The top level root node
2 (Recovery :num_retries 6 :name NavigateRecovery
3 (PipelineSequence :name NavigateWithReplanning
4 (RateController :args (hz 1)
5 (Recovery :num_retries 1 :name ComputePathToPose
6 (Action :ID ComputePathToPose :args (goal \$goal path \$path planner_id GridBased))
7 (ReactiveFallback :name ComputePathToPoseRecoveryFallback
8 (Condition :ID GoalUpdated)
9 (Action :ID ClearEntireCostmap :name ClearGlobalCostmap_Context1
10 :args (service_name global_costmap/clear_entirely_global_costmap)))))
11 (Recovery :num_retries 1 :name FollowPath
12 (Action :ID FollowPath :args (path \$path controller_id FollowPath))
13 (ReactiveFallback :name FollowPathRecoveryFallback
14 (Condition :ID GoalUpdated)
15 (Action :ID ClearEntireCostmap :name ClearLocalCostmap_Context2
16 :args (service_name local_costmap/clear_entirely_local_costmap)))))
17 (ReactiveFallback :name RecoveryFallback
18 (Condition :ID GoalUpdated)
19 (RoundRobin :name RecoveryActions
20 (Sequence :name ClearingActions
21 (Action :ID ClearEntireCostmap :name ClearLocalCostmap_Subtree3
22 :args ( service_name local_costmap/clear_entirely_local_costmap))
23 (Action :ID ClearEntireCostmap :name ClearGlobalCostmap_Subtree4
24 :args (service_name global_costmap/clear_entirely_global_costmap)))
25 (Action :ID Spin :args (spin_dist 1.57))
26 (Action :ID Wait :args (wait_duration 5))
27 (Action :ID BackUp :args (backup_dist 0.15 backup_speed 0.025)))))))
Figure 18: The graphical representation of ROS2 Nav2 BT Listing 11p11.
Appendix D The Mars rover example from Biggar and Zamani (2020); Wang et al. (2024).
Listing 12: The BTF model of the Mars rover example from Biggar and Zamani (2020); Wang et al. (2024).
1( ;example from
2 ; A framework for formal verification of behavior trees with linear temporal logic.(2020)
3 ; O. Biggar and M. Zamani.
4 ; also presented in
5 ; Enabling Behaviour Tree Verification via a Translation to BIP. (2024)
6 ; Q Wang, H Dai, Y Zhao, M Zhang, S Bliudze
7 (defsv meteo ; this defines the meteo state variable
8 :states (MInit Normal Storm) ; MInit an undefined init state
9 :init MInit ; the following transitions forbid coming back to MInit
10 :transitions ((MInit Normal) (MInit Storm) (Storm Normal) (Normal Storm)))
11
12 (defsv battery
13 :states (BInit Good Low) ; BInit just to says that we do not know
14 :init BInit ; the following transitions forbid coming back to BInit
15 :transitions ((BInit Good) (BInit Low) (Low Good) (Good Low)))
16
17 (defsv panel
18 :states (PInit Folded Unfolded) ; PInit just to says that we do not know
19 :init PInit ; the following transitions forbid coming back to PInit
20 :transitions ((PInit Folded) (PInit Unfolded) (Unfolded Folded) (Folded Unfolded)))
21
22(BehaviorTree :name mars_rover
23 (Fallback
24 (Sequence
25 (Eval (= battery Low)) ; if the battery is low
26 (Action :ID UnfoldPanels :name unfold_panels :SF) ; we unfold the panel
27 (Eval (:= panel Unfolded))) ; and the panel SV becomes unfolded
28 (Sequence
29 (Eval (= meteo Storm)) ; if the meteo is a storm
30 (Action :ID Hibernate :name hibernate :SF) ; we hibernate
31 (Eval (:= panel Folded))) ; and we fold the panel
32 (Sequence
33 (Action :ID DataReady :name dataready :SF)
34 (Action :ID Send :name send :SF)))))
35
36; prove absent ( mars_rover/3/state Unfolded and mars_rover/1/state Storm )